The length of a rectangle is twice the width. The area is 72yd^2. Find the length and width
Answer by ewatrrr(24785)

 
Hi
Let x and 2x represent the width and length respectively
Question states***
x*2x = 72yd^2
2x^2 = 72yd^2
x^2 = 36yd^2 |tossing out the negative solution for # of units
x = 6yd, the width. The length is 12yd